OpenStudy (anonymous):

(Stoichiometry) The annual production of sulfur dioxide from burning coal and fossil fuels, for auto exhaust, and from other activities is about 26 million tons. How much sulfur, present in the original materials, would result in that quantity of SO2? S+O2 --> SO2

OpenStudy (abb0t):

1. start by balancing your chemical equation (if it isn't already). 2. then converting tons \(\rightarrow\) grams. See, you're told that you start with 2.6 \(\times\) 10\(^7\) tons. \(\sf \color{red}{NOTE}\): 1 ton = 907185 grams When you complete that, let me know and we can proceed.

OpenStudy (abb0t):

\(\sf \color{blue}{HINT}\): (2.6 \(\times\)10\(^7\) tons) \(\times\) \(\sf \color{purple}{\frac{907185~gram}{1~ton}}\) = ___________ grams SO\(_2\)